# VMRNN: Integrating Vision Mamba and LSTM for Efficient and Accurate Spatiotemporal Forecasting

### v1

@@ 2.1 Convolution-based Architecture
Prior models integrating CNNs and RNNs employ various strategies to better capture spatiotemporal dependencies to improve prediction accuracy. ConvLSTM [41] extends FC-LSTM [44] by replacing fully connected operations with convolutional operations to learn spatiotemporal dependencies. PredRNN [54] proposes the Spatiotemporal LSTM (ST-LSTM) module, which simultaneously models spatiotemporal information by transferring hidden states in horizontal and vertical directions. PredRNN++ [52] designs a Gradient Highway unit to solve the vanishing gradient problem in PredRNN. E3D-LSTM [53] replaces the 2D convolution in the ST-LSTM module with 3D convolution, allowing the ST-LSTM module to remember more previous information and achieve better prediction performance. MIM [56] replaces the forget gate in the ST-LSTM module with two recurrent units to solve the non-stationary information learning in prediction. CrevNet [61] proposes a CNN-based reversible network to learn complex spatiotemporal dependencies. PhyDNet [14] introduces physical knowledge into a CNN-based model to improve prediction quality. The above models [41, 54, 52, 53, 56, 61, 14] improve their ability to capture spatiotemporal dependencies from different perspectives and achieve excellent results. However, despite their widespread usage, convolutional methods have inherent limitations in capturing spatiotemporal dependencies due to their local nature.
@@ 2.2 Transformer-based Architecture
The widespread use of Transformer [50] in the field of natural language processing (NLP) led researchers to introduce it to the vision domain. ViT [4] pioneered the direct application of the Transformer architecture to image classification and achieved great success. Swin Transformer [31] achieves outstanding performance on image classification, semantic segmentation, and object detection tasks due to its shifted window mechanism and hierarchical design. SwinLSTM [48] successfully integrates Swin Transformer [31] with LSTM which stands out as a strong spatiotemporal prediction baseline. Yet, the attention mechanism in ViT-based methods entails quadratic complexity relative to image sizes, leading to substantial computational costs, and marking a significant limitation.
@@ 3.2 VMRNN Module
ConvLSTM [41] incorporates convolutional operators into both input-to-state and state-to-state transitions, effectively addressing the limitations of FC-LSTM [44] in handling spatio-temporal data.
ConvLSTM [41]’s key equations are as follows:
where denotes the convolution operator and denotes the Hadamard product, and , , input, forget, output gate respectively.
In Eqn. 9, . Therefore, we fuse the three gates: , , into one gate, named filter gate .
We follow SwinLSTM and propose the VMRNN Module, which is detailedly presented in Fig. 3 (a). In VMRNN, the information of cell states and hidden states is updated horizontally to capture long-term and short-term temporal dependencies. Meanwhile, the VSS Blocks vertically learn spatial dependencies. The key equations of VMRNN are shown in Eqn. 12, where VSB denotes the VSS Blocks and LP denotes the Linear Projection:
@@ 3.3 VSS Block
where is four different scanning directions. and corresponding to the scan expand and scan merge operations. The selective scan space state sequential model (S6) in Eqn. 14 is the core SSM operator of the VSS block. It enables each element in a 1D array to interact with any of the previously scanned samples through a compressed hidden state. We plot the equations of S6 process in Fig. 2 (c). in S6 Block stage.
@@ 4.1 Implementations
We employ the Mean Squared Error (MSE) loss function across all three datasets. For KTH [39] and TaxiBJ [62], our methodology aligns with OpenSTL [47]. For the Moving MNIST [44] dataset, we adhere to the experimental setup detailed in [48]. The precise model parameters, hyper-parameters(including batch size, learning rate, and training epochs), and training machines utilized for each dataset are comprehensively enumerated in Table 1.

### v2

@@ 2.1 Convolution-based Architecture
Previous models that merge CNNs with RNNs adopt a range of tactics to more effectively grasp the nuances of spatiotemporal relationships, aiming to enhance predictive precision. ConvLSTM [41] evolves from FC-LSTM [44] by integrating convolutional operations instead of fully connected ones, facilitating the learning of spatiotemporal interdependencies. PredRNN [54] and its Spatiotemporal LSTM (ST-LSTM) unit mark a significant step, enabling the concurrent processing of spatiotemporal data by propagating hidden states both horizontally and vertically. Building on this, PredRNN++[52] contributes the Gradient Highway unit to mitigate the vanishing gradient issue encountered by its predecessor. Meanwhile, E3D-LSTM[53] enhances the ST-LSTM’s memory capacity by implementing 3D convolutions. The MIM model [56] reimagines the ST-LSTM’s forget gate with dual recurrent units to better address the learning of non-stationary information within predictions. CrevNet [61] employs a CNN-based reversible architecture to decode complex spatiotemporal patterns. Additionally, PhyDNet [14] embeds physical principles into CNN frameworks to refine the quality of its predictions. Collectively, these models [41, 54, 52, 53, 56, 61, 14] showcase a variety of approaches to enhance the capture of spatiotemporal dependencies and have garnered commendable outcomes. Nonetheless, conventional convolutional methodologies are constrained in their ability to seize spatiotemporal dependencies due to their intrinsic localized operation.
@@ 2.2 Transformer-based Architecture
The adoption of the Transformer model [50], initially celebrated in natural language processing, has prompted its exploration within the realm of computer vision. The Vision Transformer (ViT)[4] broke new ground by directly applying Transformer architecture to image classification, demonstrating impressive results. Further advancing this domain, the Swin Transformer[31] delivers remarkable achievements across a spectrum of tasks such as image classification, semantic segmentation, and object detection, thanks to its innovative shifted window strategy and hierarchical structure. Building on this, SwinLSTM [48] innovatively merges the Swin Transformer [31] with LSTM, establishing a new robust benchmark for spatiotemporal forecasting. However, ViT and its derivatives exhibit a notable drawback: the attention mechanism’s quadratic complexity in relation to image size, which imposes considerable computational demands.
@@ 3.2 VMRNN Module
ConvLSTM [41] incorporates convolutional operators into the information transition from both input and state and state to state, effectively addressing the limitations of FC-LSTM [44] in handling spatio-temporal data.
VMRNN Module removes all weights and biases in ConvLSTM to obtain Eqn. 3:
In Eqn. 3, . Then these three gates: , , are fused into one filter gate .
We propose the VMRNN Module, which is detailedly presented in Fig. 3 (a). In VMRNN, the long-term and short-term temporal dependencies are captured by updating the information of cell states , and hidden states is updated from a horizontal perspective. And the VSS Blocks vertically learn spatial dependencies.
We show the key equations of VMRNN in Eqn. 6, where VSB means the VSS Blocks in Sec. 3.3 and LP is short for the Linear Projection:
@@ 3.3 VSS Block
where is four different scanning directions. and corresponding to the scan expand and scan merge operations. The selective scan space state sequential model (S6) in Eqn. 8 is the core SSM operator of the VSS block. It enables each element in a 1D array to interact with any of the previously scanned samples through a compressed hidden state. We plot the equations of S6 process in Fig. 2 (c). in S6 Block stage.
@@ 4.1 Implementations
We employ the Mean Squared Error (MSE) loss function across all three datasets. For KTH [39] and TaxiBJ [62], our methodology aligns with OpenSTL [47]. For the Moving MNIST [44] dataset, we adhere to the experimental setup detailed in [48]. The precise model parameters, hyper-parameters(including batch size, learning rate, and training epochs), and training machines utilized for each dataset are comprehensively enumerated in Table 1. For TaxiBJ, we train 200 epochs with a learning rate of 4e-4 with a single A6000 GPU, using a batch size of 16. For KTH, we train 100 epochs with a learning rate of 5e-4 and 1e-4 for KTH20 and KTH40, respectively. For the Moving MNIST dataset, we adhere to the experimental setup detailed in [48] and train 2000 epochs with a learning rate of 5e-5 and a batch size of 8 using a single RTX 3090Ti GPU.